OpenStudy (lexi724):

Fan and Medal for checking my work! What is the length of the third side of the right triangle? A.15 sqrt.3 B.20 cm C. 15 sqrt.2 D.25 cm so A=15^2 B=15^2 and we have to find c so 225+225= 450 So C= 21.21

OpenStudy (lexi724):

so answer choice....... C

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

Is there a picture?

OpenStudy (igreen):

Yep, you got it.

OpenStudy (lexi724):

so it is C?

OpenStudy (lexi724):

awesome thanks!

OpenStudy (igreen):

\(a^2 + b^2 = c^2\) \(15^2 + 15^2 = c^2\) \(225 + 225 = c^2\) \(450 = c^2\) \(c = \sqrt{450}\) \(c = \sqrt{225} \times \sqrt 2\) \(c = 15 \times \sqrt 2\) \(c = 15 \sqrt 2\)

OpenStudy (igreen):

\(\Huge\color{lime} \checkmark\)

OpenStudy (lexi724):

thanks guys that was the last one ofr pythagoreum theorum---- next i have geometric mean and im no good at that

OpenStudy (igreen):

Lol

OpenStudy (lexi724):

i have some random ones before that so no worries